Create a simplified expression for the perimeter of this rectangle. Horizontally it is labelled 3x-11, and vertically it is labeled x+9. What do I do??


1
Expert's answer
2022-02-09T17:54:00-0500

Here the length of the rectangle is "L=3x-11" and the width is "W= x+9". To find the perimeter "(P)"  of this rectangle "P=2(L+W)=2((3x-11)+(x+9))=2(4x-2)=8x-4"  

The perimeter of this rectangle is "P=(8x-4)"